OpenStudy (anonymous):

GIVING MEDALS!!! Write the equation of a line in point-slope form that is parallel to the line y=3x+6 and passes through the point (-5,8)

OpenStudy (anonymous):

remember that if its parallel the slope is the same and just put that point in point slope with that slope.

OpenStudy (anonymous):

So the answer is y=3x+23? @lonnie455rich

OpenStudy (anonymous):

that is in slope intercept form. didn't it ask for point slope which is y-y=m(x-x)

OpenStudy (anonymous):

Oh yeah! Thanks :D

OpenStudy (anonymous):

y-8=3(x-(-5))? @lonnie455rich

OpenStudy (anonymous):

it would be x+5 but ya. remember when you have 2 negatives like that it makes it positive.